OpenStudy (anonymous):

Assume that m>0. Solve for v (eliminating m), being careful to choose a solution satisfying the assumption: v2 = 9 + 64 m v = 3 + 8 m

OpenStudy (anonymous):

is that v^2 or 2v

OpenStudy (anonymous):

v^2

OpenStudy (anonymous):

so I would solve for m in the v=3+8m so what would you get

OpenStudy (anonymous):

i have no idea

OpenStudy (anonymous):

solve for m v=3+8m need to get m by itself substract 3 from both sides of equal sign v-3=3-3+8m rewrite as 8m=v-3 divdie by 8 m=1/8 v -3/8 now sub this into the first equation v^2=9+64(1/8 v-3/8) you try to finish this let me know what you get

OpenStudy (anonymous):

-18.25=v

OpenStudy (anonymous):

???

OpenStudy (anonymous):

is that one eight ? in the m= equation?

OpenStudy (anonymous):

v^2=9+64(1/8 v-3/8) v^2=9+64/8 v-64(3)/8 v^2=9+8v-8(3) v^2=9+8v-24 v^2-8x+15=0 (v-5)(v-3)=0 so v=5 and v=3 5=3+8m 8m=2 m=2/8=1/4 then use v=3 3=3+8m 8m=0 m=0 so solution is v=5 and m= 1/4